# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ IV-O3 Monte Carlo Tree Search (OmegaPRM)
Luo et al. [26] addresses the persistent challenge of enhancing multi-step reasoning—particularly in mathematical problem solving—by introducing an automated process supervision framework that eschews human annotation. Conventional Outcome Reward Models (ORMs) often fall short in tasks with lengthy or multi-hop reasoning chains because they do not adequately reward or penalize intermediate steps. To overcome this, the authors propose OmegaPRM, a novel divide-and-conquer style Monte Carlo Tree Search (MCTS) algorithm that efficiently collects high-quality process supervision data. OmegaPRM employs a binary search strategy to swiftly locate the first error in a chain-of-thought (CoT), thereby balancing positive and negative examples and significantly reducing computational overhead. This method enabled the generation of over 1.5 million supervision annotations, which were used to train a Process Reward Model (PRM) via pointwise soft label training objectives based on Monte Carlo estimations. When integrated with a weighted self-consistency algorithm, the approach improved the math reasoning performance of the instruction-tuned Gemini Pro model, boosting its success rate on the MATH benchmark from 51% to 69.4%—a 36% relative improvement. Importantly, this fully automated framework offers a cost-effective alternative to human annotation while demonstrating superior performance compared to previous methods such as PRM800K and Math-Shepherd, marking a significant step forward in developing more reliable and efficient process supervision techniques for LLM reasoning tasks.
IV-O4 Self-taught Reasoner with Tools
@@
Li et al. [83] introduce START (Self-Taught Reasoner with Tools), an innovative framework that enhances the reasoning capabilities of large language models by integrating external computational tools into the long chain-of-thought (CoT) process. Traditional large reasoning models, such as OpenAI-o1 and DeepSeek-R1, exhibit strong reasoning abilities but are often hindered by hallucinations and inefficiencies stemming from their exclusive reliance on internal reasoning. START addresses these limitations by incorporating a self-learning framework that enables the model to perform complex computations, self-check, explore alternative methods, and self-debug through tool calls like a Python interpreter. The core innovations include the Hint-infer mechanism, which strategically injects context-specific hints (e.g., “Wait, maybe using Python here is a good idea”) to activate latent tool-using abilities without needing demonstration data, and the Hint Rejection Sampling Fine-Tuning (Hint-RFT) process, which refines reasoning trajectories via scoring, filtering, and fine-tuning. Applying to the QwQ-32B model, START demonstrates substantial improvements across various benchmarks—including AMC, AIME, and coding challenges—achieving competitive performance that rivals state-of-the-art models while significantly reducing debug errors and enhancing overall problem-solving accuracy.
IV-P Test-Time Compute scaling
@@
Brown et al. [84] explore an often-overlooked axis of scaling—namely, inference compute—by systematically examining how repeated sampling can enhance the performance of language models. Traditionally, LLMs make a single attempt at generating an answer, but this study demonstrates that generating multiple, diverse candidate solutions and selecting the best one can significantly improve the fraction of problems solved, a metric referred to as coverage. Empirical results indicate that coverage scales in a log-linear fashion with the number of samples, following an exponentiated power law. In domains such as coding and formal proofs—where automatic verifiers like unit tests are available—this approach directly translates into measurable performance gains. For instance, on the SWE-bench Lite benchmark, the DeepSeek-Coder-V2-Instruct model’s performance increases from 15.9% to 56% when the number of samples is raised from one to 250, outstripping the single-sample state-of-the-art of 43%. However, in domains lacking automatic verification, traditional selection strategies like majority voting and reward model scoring tend to plateau after a few hundred samples. The implementation strategy involves generating a high number of independent responses with elevated sampling temperatures for diversity, followed by selecting the best candidate based on established criteria.
@@
Huggingface team [85] investigated dynamic inference strategies—such as repeated sampling methods—that allow models to “think longer” on challenging tasks like complex math problems. They systematically evaluate several search-based techniques including majority voting (self-consistency), Best-of-N sampling, beam search augmented with process reward models (PRMs), and a novel Diverse Verifier Tree Search (DVTS) designed to enhance candidate diversity under high compute budgets. Their experimental pipeline involves generating multiple candidate solutions, scoring intermediate reasoning steps with a PRM, and selecting optimal outputs based on these scores. Remarkably, the study demonstrates that smaller models (e.g., Llama 3.2 1B) can, with sufficient test-time compute, rival or even surpass the performance of larger counterparts (such as Llama 3.1 70B) on the MATH-500 benchmark, achieving up to 55% accuracy. The findings also reveal that the effectiveness of each search strategy varies with problem difficulty—beam search tends to excel on harder problems while Best-of-N sampling is better suited for simpler ones—and that an adaptive, compute-optimal scaling approach can yield significant performance gains.
IV-Q SageAttention
@@
SageAttention [122] is a novel quantization method specifically designed to accelerate the attention mechanism in transformers—a critical component that traditionally incurs an O(N²) computational cost with long sequence lengths. While existing quantization techniques predominantly optimize the linear layers, SageAttention directly targets attention by employing a 4/8-bit quantization for the Q and K matrices and FP8/16 for the P and V matrices, complemented by smoothing methods for Q and V. It dynamically adjusts quantization parameters across timesteps and layers, serving as a drop-in replacement for PyTorch’s scaled_dot_product_attention. Experimental results are compelling: SageAttention achieves approximately a 3× speedup over FlashAttention2, outperforming FlashAttention2 and xformers by about 2.1× and 2.7× in operations per second, respectively, while maintaining roughly 99% of the original model performance. Moreover, extensive evaluations demonstrate that this approach results in negligible end-to-end performance loss across a range of applications, including large language processing, image generation, and video generation.
The standard self-attention is computed as follows [123]:
@@
IV-R Speculative Knowledge Distillation (SKD)
@@
Google introduced Speculative Knowledge Distillation (SKD) [86], a novel approach designed to overcome the inherent limitations of conventional knowledge distillation methods—namely, supervised KD and on-policy KD—in bridging the gap between teacher and student models. Traditional supervised KD suffers from a distribution mismatch between static training datasets and the dynamic outputs produced during inference, while on-policy KD often relies on lower-quality, student-generated samples that can yield inaccurate teacher feedback. In contrast, SKD leverages an interleaved sampling strategy in which the student model initially generates a draft sequence of tokens, which are then evaluated against the teacher model’s top-K predictions. If a student-proposed token ranks within the top 25 of the teacher’s probability distribution, it is retained; otherwise, it is replaced by the teacher’s highest probability token. This adaptive token replacement mechanism allows SKD to generate high-quality training data on-the-fly that is better aligned with the student’s inference-time distribution. Evaluated across diverse text generation tasks—including translation, summarization, mathematical problem-solving, and instruction following—SKD consistently outperforms supervised and on-policy KD methods, enhancing speculative decoding performance in the distilled model.
IV-S Self-Taught Reasoner (STaR)
@@
Zelikman et al. [17] proposed the Self-Taught Reasoner (STaR), an innovative iterative training technique that enables language models to improve their reasoning capabilities by learning from their own generated rationales. Traditional chain-of-thought prompting requires either extensive datasets of annotated rationales or relies on few-shot inference, which can compromise accuracy. STaR circumvents these limitations through a simple yet effective loop: using a small set of few-shot examples with rationales, the model generates step-by-step explanations (rationales) and answers for a large collection of problems. When the generated answer is incorrect, the model is prompted again—this time with a hint that includes the correct answer—to produce a revised rationale. The model is then fine-tuned on both the initially correct rationales and these corrected examples. Repeating this process over multiple iterations (typically 30–40) enables the model to refine its reasoning skills progressively. Experimental results show that STaR significantly improves performance on various reasoning tasks, such as GSM8K and CommonsenseQA, achieving gains comparable to those from fine-tuning a model 30 times larger. Notably, on GPT-J, STaR boosts performance from 5.8% to 10.7%, demonstrating its potential to reduce the reliance on large, manually labeled rationale datasets while effectively enabling models to learn from their mistakes.
@@
V-STaR [81] represents a significant advancement in self-improvement for large language models by harnessing the overlooked value of incorrect solutions. Traditional approaches like STaR focus solely on fine-tuning models using correct self-generated solutions, discarding a wealth of information present in the erroneous outputs. In contrast, V-STaR employs a novel strategy where both correct and incorrect solutions are utilized to form preference pairs, which are then used to train a DPO-based verifier. This verifier evaluates and ranks candidate solutions during inference, selecting the most promising output. By iteratively refining both the generator and the verifier, V-STaR achieves an impressive 4% to 17% improvement in test accuracy on common code generation and math reasoning benchmarks with LLaMA2 models, outperforming conventional self-consistency and majority voting approaches, particularly for smaller candidate sets.
@@
Quiet-STaR [18] is a novel generalization of the Self-Taught Reasoner (STaR) approach that teaches language models to ”think before they speak” by generating token-level rationales—essentially internal thoughts—alongside each token prediction. Instead of being limited to generating final answers, Quiet-STaR trains models to output parallel sequences of ”thoughts” (delimited by learnable ¡—startofthought—¿ and ¡—endofthought—¿ tokens) for each token using a tokenwise parallel sampling algorithm with a specially designed attention mask. A mixing head then integrates these thought outputs to inform next-token predictions, while REINFORCE is used to optimize the generated thoughts based on how much they improve the log-likelihood of future tokens. Iteratively training on large diverse corpora (e.g., OpenWebMath, C4) leads to significant zero-shot performance gains on reasoning-heavy tasks—improving GSM8K scores from 5.9% to 10.9% and CommonsenseQA from 36.3% to 47.2% for a Mistral 7B model—demonstrating that embedding reflective reasoning into every token can help models more effectively predict complex, unstated logic without additional fine-tuning.
IV-T Self-Taught Evaluators
@@
Meta team [82] proposes a novel iterative self-improvement framework for model-based evaluation that entirely forgoes the need for human-labeled preference data. Traditionally, training evaluators—whether for reward modeling in reinforcement learning or as stand-ins for human evaluators—relies on extensive collections of human judgments, a process that is both costly and susceptible to obsolescence as models evolve. Instead, the authors introduce a self-taught evaluator approach where an LLM is trained using synthetic data generated from unlabeled instructions. In each iteration, the model produces contrasting outputs—one high-quality and one intentionally sub-optimal—followed by generating detailed reasoning traces and final judgments for these response pairs. These synthetic preference pairs are then used to iteratively fine-tune the evaluator, with each cycle leveraging the improved judgments from the previous iteration. Empirical results on RewardBench demonstrate a marked improvement: the evaluator based on Llama3-70B-Instruct sees its performance increase from 75.4% to 88.3% (or 88.7% with majority voting), matching the effectiveness of top reward models trained on human annotations. This synthetic, iterative strategy not only reduces reliance on costly human data but also allows for customizable evaluation criteria, though it also highlights potential challenges such as the risk of amplifying initial model biases over successive iterations.
IV-U Self-Enhanced Test-Time Scaling (SETS)
@@
Google team presents Self-Enhanced Test-Time Scaling (SETS) [124], a novel framework that leverages the self-verification and self-correction capabilities of advanced LLMs to overcome the diminishing returns observed in conventional test-time scaling methods. Traditional approaches, such as repeated sampling with majority voting or employing task-specific reward models, have struggled to yield proportional performance gains as computational resources increase. In contrast, SETS integrates a three-step iterative loop—generating multiple solution candidates, self-verifying these outputs using the model’s inherent reasoning abilities, and then applying targeted self-correction to refine erroneous responses—into a unified process. By incorporating majority voting over these refined outputs, the framework enhances accuracy and demonstrates more favorable scaling characteristics. Experimental evaluations on challenging planning and reasoning benchmarks indicate that SETS can achieve up to an 8.7% improvement in accuracy while reducing compute requirements by 22% compared to traditional sampling-heavy baselines. This approach underscores a promising direction for exploiting the emergent self-refinement abilities of LLMs, thereby pushing the boundaries of test-time computation in complex reasoning tasks.
IV-V Stream of Search (SoS)
@@
Gandhi et al. [125] introduced the ”Stream of Search” (SoS) approach, a training methodology designed to teach language models to perform structured search and backtracking by representing the entire search process as a continuous sequence—a flattened string of actions. Unlike conventional training, which exposes models only to optimal reasoning trajectories, SoS incorporates all intermediate steps, including mistakes and corrective backtracking, thereby mimicking human problem-solving behavior. The method is demonstrated on the challenging Countdown game, where a set of input numbers must be combined through arithmetic operations to reach a target number. Using a heuristic problem solver, the authors generated 500,000 chain-of-thought/search examples; notably, only 57% of these samples led to correct solutions, highlighting the inherent complexity of the search process. Training a transformer-based language model on this rich, diverse dataset improved search accuracy by 25% compared to models trained solely on optimal paths. Furthermore, by applying policy improvement techniques such as Self-Taught Reasoner (STaR) and Advantage-Induced Policy Alignment (APA), the fine-tuned models achieved an additional  6% gain, ultimately solving 36% of previously unsolvable problems. Overall, the study demonstrates that exposing language models to the full spectrum of search behaviors—including errors—enables them to learn more robustly and self-correct, thereby improving their performance on complex reasoning tasks.
IV-W Reflective Augmentation (RefAug)

### v2

@@ VI-O3 Monte Carlo Tree Search (OmegaPRM)
Luo et al. [luo2024improve] addresses the persistent challenge of enhancing multi-step reasoning particularly in mathematical problem solving by introducing an automated process supervision framework that eschews human annotation. Conventional Outcome Reward Models (ORMs) often fall short in tasks with lengthy or multi-hop reasoning chains because they do not adequately reward or penalize intermediate steps. To overcome this, the authors propose OmegaPRM, a novel divide-and-conquer style Monte Carlo Tree Search (MCTS) algorithm that efficiently collects high-quality process supervision data. OmegaPRM employs a binary search strategy to swiftly locate the first error in a chain-of-thought (CoT), thereby balancing positive and negative examples and significantly reducing computational overhead. This method enabled the generation of over 1.5 million supervision annotations, which were used to train a Process Reward Model (PRM) via pointwise soft label training objectives based on Monte Carlo estimations. When integrated with a weighted self-consistency algorithm, the approach improved the math reasoning performance of the instruction-tuned Gemini Pro model, boosting its success rate on the MATH benchmark from 51% to 69.4% a 36% relative improvement. Importantly, this fully automated framework offers a cost-effective alternative to human annotation while demonstrating superior performance compared to previous methods such as PRM800K and Math-Shepherd, marking a significant step forward in developing more reliable and efficient process supervision techniques for LLM reasoning tasks.
VI-O4 Self-taught Reasoner with Tools
@@
Li et al. [li2025start] introduce START (Self-Taught Reasoner with Tools), an innovative framework that enhances the reasoning capabilities of large language models by integrating external computational tools into the long chain-of-thought (CoT) process. Traditional large reasoning models, such as OpenAI-o1 and DeepSeek-R1, exhibit strong reasoning abilities but are often hindered by hallucinations and inefficiencies stemming from their exclusive reliance on internal reasoning. START addresses these limitations by incorporating a self-learning framework that enables the model to perform complex computations, self-check, explore alternative methods, and self-debug through tool calls like a Python interpreter. The core innovations include the Hint-infer mechanism, which strategically injects context-specific hints (e.g., “Wait, maybe using Python here is a good idea”) to activate latent tool-using abilities without needing demonstration data, and the Hint Rejection Sampling Fine-Tuning (Hint-RFT) process, which refines reasoning trajectories via scoring, filtering, and fine-tuning. Applying to the QwQ-32B model, START demonstrates substantial improvements across various benchmarks including AMC, AIME, and coding challenges achieving competitive performance that rivals state-of-the-art models while significantly reducing debug errors and enhancing overall problem-solving accuracy.
VI-P Test-Time Compute scaling
@@
Brown et al. [brown2024large] explore an often-overlooked axis of scaling namely, inference compute by systematically examining how repeated sampling can enhance the performance of language models. Traditionally, LLMs make a single attempt at generating an answer, but this study demonstrates that generating multiple, diverse candidate solutions and selecting the best one can significantly improve the fraction of problems solved, a metric referred to as coverage. Empirical results indicate that coverage scales in a log-linear fashion with the number of samples, following an exponentiated power law. In domains such as coding and formal proofs where automatic verifiers like unit tests are available this approach directly translates into measurable performance gains. For instance, on the SWE-bench Lite benchmark, the DeepSeek-Coder-V2-Instruct model’s performance increases from 15.9% to 56% when the number of samples is raised from one to 250, outstripping the single-sample state-of-the-art of 43%. However, in domains lacking automatic verification, traditional selection strategies like majority voting and reward model scoring tend to plateau after a few hundred samples. The implementation strategy involves generating a high number of independent responses with elevated sampling temperatures for diversity, followed by selecting the best candidate based on established criteria.
@@
Huggingface team [huggingface2023scaling] investigated dynamic inference strategies such as repeated sampling methods that allow models to “think longer” on challenging tasks like complex math problems. They systematically evaluate several search-based techniques including majority voting (self-consistency), Best-of-N sampling, beam search augmented with process reward models (PRMs), and a novel Diverse Verifier Tree Search (DVTS) designed to enhance candidate diversity under high compute budgets. Their experimental pipeline involves generating multiple candidate solutions, scoring intermediate reasoning steps with a PRM, and selecting optimal outputs based on these scores. Remarkably, the study demonstrates that smaller models (e.g., Llama 3.2 1B) can, with sufficient test-time compute, rival or even surpass the performance of larger counterparts (such as Llama 3.1 70B) on the MATH-500 benchmark, achieving up to 55% accuracy. The findings also reveal that the effectiveness of each search strategy varies with problem difficulty beam search tends to excel on harder problems while Best-of-N sampling is better suited for simpler ones and that an adaptive, compute-optimal scaling approach can yield significant performance gains.
VI-Q SageAttention
@@
SageAttention [zhang2024sageattention] is a novel quantization method specifically designed to accelerate the attention mechanism in transformers a critical component that traditionally incurs an O(N²) computational cost with long sequence lengths. While existing quantization techniques predominantly optimize the linear layers, SageAttention directly targets attention by employing a 4/8-bit quantization for the Q and K matrices and FP8/16 for the P and V matrices, complemented by smoothing methods for Q and V. It dynamically adjusts quantization parameters across timesteps and layers, serving as a drop-in replacement for PyTorch’s scaled_dot_product_attention. Experimental results are compelling: SageAttention achieves approximately a 3× speedup over FlashAttention2, outperforming FlashAttention2 and xformers by about 2.1× and 2.7× in operations per second, respectively, while maintaining roughly 99% of the original model performance. Moreover, extensive evaluations demonstrate that this approach results in negligible end-to-end performance loss across a range of applications, including large language processing, image generation, and video generation.
The standard self-attention is computed as follows [vaswani2017attention]:
@@
VI-R Speculative Knowledge Distillation (SKD)
@@
Google introduced Speculative Knowledge Distillation (SKD) [xu2024speculative], a novel approach designed to overcome the inherent limitations of conventional knowledge distillation methods namely, supervised KD and on-policy KD in bridging the gap between teacher and student models. Traditional supervised KD suffers from a distribution mismatch between static training datasets and the dynamic outputs produced during inference, while on-policy KD often relies on lower-quality, student-generated samples that can yield inaccurate teacher feedback. In contrast, SKD leverages an interleaved sampling strategy in which the student model initially generates a draft sequence of tokens, which are then evaluated against the teacher model’s top-K predictions. If a student-proposed token ranks within the top 25 of the teacher’s probability distribution, it is retained; otherwise, it is replaced by the teacher’s highest probability token. This adaptive token replacement mechanism allows SKD to generate high-quality training data on-the-fly that is better aligned with the student’s inference-time distribution. Evaluated across diverse text generation tasks including translation, summarization, mathematical problem-solving, and instruction following SKD consistently outperforms supervised and on-policy KD methods, enhancing speculative decoding performance in the distilled model.
VI-S Self-Taught Reasoner (STaR)
@@
Zelikman et al. [zelikman2022star] proposed the Self-Taught Reasoner (STaR), an innovative iterative training technique that enables language models to improve their reasoning capabilities by learning from their own generated rationales. Traditional chain-of-thought prompting requires either extensive datasets of annotated rationales or relies on few-shot inference, which can compromise accuracy. STaR circumvents these limitations through a simple yet effective loop: using a small set of few-shot examples with rationales, the model generates step-by-step explanations (rationales) and answers for a large collection of problems. When the generated answer is incorrect, the model is prompted again this time with a hint that includes the correct answer to produce a revised rationale. The model is then fine-tuned on both the initially correct rationales and these corrected examples. Repeating this process over multiple iterations (typically 30–40) enables the model to refine its reasoning skills progressively. Experimental results show that STaR significantly improves performance on various reasoning tasks, such as GSM8K and CommonsenseQA, achieving gains comparable to those from fine-tuning a model 30 times larger. Notably, on GPT-J, STaR boosts performance from 5.8% to 10.7%, demonstrating its potential to reduce the reliance on large, manually labeled rationale datasets while effectively enabling models to learn from their mistakes.
@@
V-STaR [hosseini2024v] represents a significant advancement in self-improvement for large language models by harnessing the overlooked value of incorrect solutions. Traditional approaches like STaR focus solely on fine-tuning models using correct self-generated solutions, discarding a wealth of information present in the erroneous outputs. In contrast, V-STaR employs a novel strategy where both correct and incorrect solutions are utilized to form preference pairs, which are then used to train a DPO-based verifier. This verifier evaluates and ranks candidate solutions during inference, selecting the most promising output. By iteratively refining both the generator and the verifier, V-STaR achieves an impressive 4% to 17% improvement in test accuracy on common code generation and math reasoning benchmarks with LLaMA2 models, outperforming conventional self-consistency and majority voting approaches, particularly for smaller candidate sets.
@@
Quiet-STaR [zelikman2024quiet] is a novel generalization of the Self-Taught Reasoner (STaR) approach that teaches language models to ”think before they speak” by generating token-level rationales essentially internal thoughts alongside each token prediction. Instead of being limited to generating final answers, Quiet-STaR trains models to output parallel sequences of ”thoughts” (delimited by learnable <|startofthought|><|startofthought|> and <|endofthought|><|endofthought|> tokens) for each token using a tokenwise parallel sampling algorithm with a specially designed attention mask. A mixing head then integrates these thought outputs to inform next-token predictions, while REINFORCE is used to optimize the generated thoughts based on how much they improve the log-likelihood of future tokens. Iteratively training on large diverse corpora (e.g., OpenWebMath, C4) leads to significant zero-shot performance gains on reasoning-heavy tasks improving GSM8K scores from 5.9% to 10.9% and CommonsenseQA from 36.3% to 47.2% for a Mistral 7B model demonstrating that embedding reflective reasoning into every token can help models more effectively predict complex, unstated logic without additional fine-tuning.
VI-T Self-Taught Evaluators
@@
Meta team [wang2024selftaughtevaluators] proposes a novel iterative self-improvement framework for model-based evaluation that entirely forgoes the need for human-labeled preference data. Traditionally, training evaluators whether for reward modeling in reinforcement learning or as stand-ins for human evaluators relies on extensive collections of human judgments, a process that is both costly and susceptible to obsolescence as models evolve. Instead, the authors introduce a self-taught evaluator approach where an LLM is trained using synthetic data generated from unlabeled instructions. In each iteration, the model produces contrasting outputs one high-quality and one intentionally sub-optimal followed by generating detailed reasoning traces and final judgments for these response pairs. These synthetic preference pairs are then used to iteratively fine-tune the evaluator, with each cycle leveraging the improved judgments from the previous iteration. Empirical results on RewardBench demonstrate a marked improvement: the evaluator based on Llama3-70B-Instruct sees its performance increase from 75.4% to 88.3% (or 88.7% with majority voting), matching the effectiveness of top reward models trained on human annotations. This synthetic, iterative strategy not only reduces reliance on costly human data but also allows for customizable evaluation criteria, though it also highlights potential challenges such as the risk of amplifying initial model biases over successive iterations.
VI-U Self-Enhanced Test-Time Scaling (SETS)
@@
Google team presents Self-Enhanced Test-Time Scaling (SETS) [chen2025setsn], a novel framework that leverages the self-verification and self-correction capabilities of advanced LLMs to overcome the diminishing returns observed in conventional test-time scaling methods. Traditional approaches, such as repeated sampling with majority voting or employing task-specific reward models, have struggled to yield proportional performance gains as computational resources increase. In contrast, SETS integrates a three-step iterative loop generating multiple solution candidates, self-verifying these outputs using the model’s inherent reasoning abilities, and then applying targeted self-correction to refine erroneous responses into a unified process. By incorporating majority voting over these refined outputs, the framework enhances accuracy and demonstrates more favorable scaling characteristics. Experimental evaluations on challenging planning and reasoning benchmarks indicate that SETS can achieve up to an 8.7% improvement in accuracy while reducing compute requirements by 22% compared to traditional sampling-heavy baselines. This approach underscores a promising direction for exploiting the emergent self-refinement abilities of LLMs, thereby pushing the boundaries of test-time computation in complex reasoning tasks.
VI-V Stream of Search (SoS)
@@
Gandhi et al. [gandhi2024stream] introduced the ”Stream of Search” (SoS) approach, a training methodology designed to teach language models to perform structured search and backtracking by representing the entire search process as a continuous sequence a flattened string of actions. Unlike conventional training, which exposes models only to optimal reasoning trajectories, SoS incorporates all intermediate steps, including mistakes and corrective backtracking, thereby mimicking human problem-solving behavior. The method is demonstrated on the challenging Countdown game, where a set of input numbers must be combined through arithmetic operations to reach a target number. Using a heuristic problem solver, the authors generated 500,000 chain-of-thought/search examples; notably, only 57% of these samples led to correct solutions, highlighting the inherent complexity of the search process. Training a transformer-based language model on this rich, diverse dataset improved search accuracy by 25% compared to models trained solely on optimal paths. Furthermore, by applying policy improvement techniques such as Self-Taught Reasoner (STaR) and Advantage-Induced Policy Alignment (APA), the fine-tuned models achieved an additional  6% gain, ultimately solving 36% of previously unsolvable problems. Overall, the study demonstrates that exposing language models to the full spectrum of search behaviors including errors enables them to learn more robustly and self-correct, thereby improving their performance on complex reasoning tasks.
VI-W Reflective Augmentation (RefAug)